please could you let me know: determine the equqtion of the line through the given point with the given slope
(2,-5) m=1
thank you

Answer by ankor@dixie-net.com(22740)

(2,-5) m=1
Use the point slope format: y - y1 = m(x-x1)
In this problem: x1=2, y1=-5, m=1
:
y - (-5) = 1(x - 2)
y + 5 = 1x - 2
y = x - 2 - 5
y = x - 7
:
:
Check this, replace x with 2
y = 2 - 7
y = -5